# Benchmarking AI Models in Software Engineering: A Review, Search Tool, and Enhancement Protocol

### v1

@@ III-B Results of the Review
In Conclusion, based on our in-depth inspection of HumanEval and MBPP and combined with the inspection of other benchmarks in our review, we obtained an overview of the limitations in current AI4SE benchmarks which we use as a guiding light for shaping our methodology proposed in section V.
IV BenchScout- Locating AI4SE Benchmarks
@@
Due to the abundance of AI4SE benchmarks, identifying the most suitable one for a specific software engineering task can be challenging. As a result, many default to evaluating their models on popular benchmarks like HumanEval [1] which has its own flaws.
@@
IV-A Context Extraction and Visualization
Approach Overview: Based on our review in section III, we compile a collection of datasets and benchmarks along with their associated papers. To present these data in a clear and understandable way, we must first map the unstructured textual content of the papers to a semistructured domain. We accomplished this using text embedding models pre-trained on relevant data, such as abstracts and academic papers.
We apply dimensionality reduction to the data to create a 2D representation that is easy for users to interpret. To further enhance understanding, we assess the similarity between benchmarks by applying clustering techniques.
Implementation: We used models from the Sentence Transformers  [^15] https://sbert.net/ library [171] to create embeddings, followed by t-SNE and k-means clustering from scikit-learn [172].
A randomized search was conducted across parameters like the embedding model, text templates, and number of clusters. We selected nine top-performing models from the Sentence Transformers leaderboard: all-MiniLM-L6-v20, facebook-dpr-ctx_encoder-multiset-base, all-mpnet-base-v2, all-distilroberta-v1, all-MiniLM-L12-v2, paraphrase-multilingual-mpnet-base-v2, paraphrase-albert-small-v2, paraphrase-multilingual-MiniLM-L12-v2, and paraphrase-MiniLM-L3-v2.
We used several text templates such as “Title Only”, “Abstract Only”, and combinations of title, abstract, and venue with and without explicit labels. The number of clusters varied from 8 to 15. Each combination of model and template was evaluated 100 times using different random seeds, and the Silhouette Score was used to assess the quality of the cluster.
The top Silhouette Score was obtained by the Title + Abstract Without Indication with the all-mpnet-base-v2 model, initially forming 8 clusters. These were subsequently reviewed and refined manually, resulting in the final 11 clusters. We developed an interactive interface around this clustering to explore and verify the clusters, naming each for easy navigation.
IV-B Additional Features
To improve the search experience, we’ve added key features for finding relevant articles. First, using a text-based search interface, users can search articles by title and abstract with fuzzy search from Fuse.js library [^16] https://www.fusejs.io/, enhancing flexibility for imprecise queries. We also added a Paper Content Tooltip for exploring a 2-D map of papers. Hovering over a point reveals a brief overview, and double-clicking leads to the paper’s DOI page for easy access. Clicking a point with Related Papers feature gives four insights: visual most similar papers (cosine similarity) in a vector space, a detailed overview of the paper with metadata, a list of papers that cite the selected paper, and a Paper Citations Graph showing connections between citing papers, illustrating research impact and relations. In Figure 3 we provide a screenshot of the first page of the tool and the current clustering of papers.
Figure 3: Screenshot of BenchScout
IV-C User Study
@@
IV-C1 Questionnaire Design
TABLE XVI: Questionnaire Design Overview
@@ IV-C1 Questionnaire Design
Participant Background - What is your professional background? - What is your role? - How familiar are you with AI4SE benchmarks? - How many years of experience do you have in this field? 5-point Likert Scale (1: Not familiar, 5: Very familiar) Experience question ($<$1, 1-3, 3-5, 5+ years)
Search Functionality - How easy was it to navigate the interface? - How intuitive was the search functionality? - How effective was the tool in finding benchmarks? - Was the visual evaluation of datasets useful? 5-point Likert Scale (1: Not useful, 5: Extremely useful) N/A
Cross- referencing Feature - How useful was the cross-referencing feature? - Did the tool help in understanding relationships between benchmarks? - Was the visual interface for benchmark similarity useful? 5-point Likert Scale, Open-ended Includes qualitative feedback option
User Experience & Feedback - How would you rate the overall user experience? - How likely are you to use the tool in your work? - Did you experience any issues or challenges? - What other tools do you use for searching benchmarks? - How does this tool compare to others? - How well does the tool meet the needs of professionals in AI4SE? Likert Scale, Open-ended Open-ended questions for in-depth feedback
The questionnaire designed for the study was divided into four key sections to gather feedback about the tool, namely the participant’s background, the quality of the search functionality, the quality of the cross-referencing feature, and the overall user experience. Table XVI is an overview of the posed questions.
IV-C2 Results and Analysis
@@
The respondents of the questionnaire were generally familiar with AI4SE, with an average familiarity rating of 3.8. There were varying levels of experience in the field, with the range between 1-3 years being most common (eight people).
@@ V BenchFrame
The HumanEval benchmark has long been the de facto standard for evaluating the code-generation capabilities of AI4SE models. It has recently been used to evaluate the latest and greatest LLMs from large companies such as Google Gemini [2] and OpenAI GPT4 [3]. Despite the great fame, however, our review in section III points towards the existence of numerous notable problems with this foundational benchmark, namely, the existence of incorrect tests, suboptimal canonical solutions, and imprecise problem definitions amongst many others.
V-A Approach
Figure 4: Outline of the general approach behind BenchFrame through a case-study of HumanEval.
To improve the quality of the given benchmark, we pursue the following approach – also illustrated in Figure 4. First, we initiate a comprehensive code review, leading to standardized observations (subsubsection V-A1). Then, we address these identified issues through a series of modifications (subsubsection V-A2), followed by a peer review (subsubsection V-A3) to ensure accuracy and reliability. Finally, we experiment with the revised benchmark to evaluate and discuss the results. Below, more details are provided for specific steps in this approach.
V-A1 Standardized Observations in Current HumanEval Benchmarks
Upon examining and manually reconstructing canonical solutions and experimenting with various HumanEval benchmark test suites, we identified several recurring issues. The system frequently produces incorrect and sub-optimal code, as canonical solutions are inefficient and fail to address critical assumptions outlined in the problem descriptions. Additionally, these solutions often lack type annotations, further complicating the evaluation process. Another significant problem is the absence of quality testing. Test suites tend to overlap with example tests from the prompt, allowing incorrect canonical solutions to pass. Moreover, there are instances where the expected outputs in the test suites do not align with the canonical solutions’ actual performance.
@@ V-A2 Modifications in HumanEvalNext
In HumanEvalNext, we address all the above issues by manually modifying all problems in the original HumanEval benchmark. We decide to improve the original HumanEval as (1) flaws in the original version persist even in improved versions such as HumanEvalPlus, and (2) the fact that HumanEval is still widely used in new literature. While in our we have gone for the original benchmark, due to the adaptable nature of BenchFrame one could opt to have an improved version of the dataset as the starting point. To summarize, these are the general changes made in HumanEvalNext and their benefits: In this work, several key improvements have been made to address the shortcomings of previous benchmarks. First, all suboptimal and incorrect canonical solutions have been fixed, which were previously missed due to insufficient testing and a lack of comprehensive quality review. Furthermore, type annotations have been added to all problems, offering valuable context and simplifying the translation to other programming languages. The original HumanEval benchmark only included type annotations for the first 30 problems, representing merely 18% of the total set of 164 problems. In addition, better support for language conversion frameworks has been incorporated. For instance, HumanEvalNext now features improved compatibility with frameworks like MultiPL-E [6], which supports translation to 18 additional programming languages by standardizing all tests to equality assertions where feasible. This change has reduced the number of incompatible problems by a factor of ten.
@@
Besides these adjustments, challenging scenarios (such as negative values, zero instances, empty inputs, and non-alphanumeric symbols) are incorporated into each task to guarantee that only top-quality AI4SE models, adept at addressing diverse situations, succeed. Accordingly, assertions are implemented within the code wherever constraints are detailed in the problem description. This measure prevents models from ignoring crucial details, thus improving the benchmark’s worth. In particular, we employ specification-based testing to assess functions; using boundary analysis, we explore combinations of within, on, and outside points. Furthermore, the test examples in the problem descriptions have been refined, which significantly affects model performance [14]. Problems with excessive test examples now feature a reduced set, distributing the evaluation workload more fairly. Lastly, spelling errors have been corrected, descriptions have been consistently formatted, and problem descriptions have been aligned with the implementations, while still leaving room for models to demonstrate intuitive problem-solving skills expected of high-quality AI4SE models. The difficulty level has also been raised by incorporating more edge cases and modifying various problems, aiming to better reflect real-world challenges faced by engineers and to mitigate issues related to data leakage and saturated performance on the HumanEval leaderboards.
To illustrate the modifications of the tests in HumanEvalNext, consider Table XVII.
@@
TABLE XVII: Comparison of test statistics between HumanEval191919based on human-eval-v2-20210705.json.
Metric HumanEval HumanEvalNext $\Delta$
Total assertions 1325 2551 ×1.92$\times$ 1.92× 1.92
Average assertions 8 16 ×2$\times$ 2× 2
Median assertions 7 11 ×1.57$\times$ 1.57× 1.57
Min. assertions 1 4 +3+3+ 3
$<$5 assertions 34 2 -94%
V-A3 Peer Review Process of HumanEvalNext
@@
To ensure the accuracy and reliability of the initial version of HumanEvalNext, an independent reviewer verified all changes. This thorough review involved verifying the clarity and completeness of the problem docstrings, checking for consistency between the problem descriptions and the canonical solutions, and ensuring that both the solutions and test cases were correct and efficient. Where inefficiencies were identified, suggestions for optimization were provided. The review also scrutinized the test cases to ensure comprehensive coverage, identifying any gaps that could allow incorrect solutions to pass.
While the initial creation of the benchmark took over 100 hours, the independent peer-review process required an additional 16 hours. As a result of this review, 1% of the problems were redesigned due to structural issues, 9% received additional test cases, and 15% underwent minor grammar or clarity improvements. All suggested changes were documented and reviewed by the original author, with every recommendation either implemented or refined further upon discussion.
Since the peer-review involved modifications beyond the test suites, completions for all models were re-run. Despite these changes, the results remained largely consistent with the original, with 40% of models showing no change in their pass@1 scores, 50% showing a 1-2% absolute change, and only 10%—previously top performers—experiencing a 5% drop. This demonstrated that the peer-review process upheld the benchmark’s robustness while refining its quality.
V-B Experimental Setup
@@
To assess the impact of the modifications applied to the benchmark, we examine the pass@1 performance of ten state-of-the-art open-source software code models using the original HumanEval alongside two enhanced variants, HumanEvalNext and EvalPlus [10]. We selected the top-performing models from the big code LLM leaderboards at the evaluation’s start: “NTQAI/Nxcode-CQ-7B-orpo”, “Qwen/CodeQwen1.5-7B”, “deepseek-ai/deepseek-coder-6.7b-instruct”, “TechxGenus/starcoder2-15b-instruct”, “ise-uiuc/Magicoder-S-DS-6.7B”, “Artigenz/Artigenz-Coder-DS-6.7B”, “HuggingFaceH4/starchat2-15b-v0.1”, “google/codegemma-7b-it”, “codeLlama/CodeLlama-13b-Instruct-hf”, and “Stabilityai/stable-code-3b”. For each task in the benchmark (164 total), the LLM is prompted using an instructional preamble asking the model to finish the implementation of the function requested in addition to providing the imports, function header, and function description with each request.
We run the inference for the models on a cluster with one NVIDIA A100 80GB GPU and 32 CPU cores. During test execution, a timeout limit of 15 seconds is utilized per function call to disregard completions that are potentially looping forever or are considered overly inefficient with regard to the canonical solutions. Each test is executed in an evaluation suite using the precautions deployed by OpenAI.
V-C Results

### v2

@@ 3.3. Limitations of Existing Benchmarks
In Conclusion, based on our in-depth inspection of HumanEval and MBPP and combined with the inspection of other benchmarks in our review, we obtained an overview of the limitations in current AI4SE benchmarks which we use as a guide in shaping our methodology proposed in Section 5.
4. BenchScout- Locating AI4SE Benchmarks
@@
Due to the abundance of AI4SE benchmarks, identifying the most suitable one for a specific software engineering task can be challenging. As a result, many default to evaluating their models on popular benchmarks like HumanEval (Chen et al., 2021) which has its own flaws.
@@
4.1. Context Extraction and Visualization
4.1.1. Overview of the Semi-Automated Pipeline.
To effectively contextualize and visualize the growing corpus of AI4SE benchmarks, we developed a semi-automated pipeline that extracts, embeds, clusters, and visualizes benchmark metadata. As illustrated in Figure 5, the system is organized into four modular layers: data collection, metadata enrichment, clustering and labeling, and front-end interaction.
The pipeline begins with benchmark sources collected from a curated spreadsheet, user-submitted suggestions, and an automated paper discovery script using the Semantic Scholar API to make suggestions. All inputs are aggregated into a processing queue, where metadata is enriched using both automated API calls and basic scraping techniques (e.g., fetching GitHub or Hugging Face README files). These metadata are transformed into dense vector representations using OpenAI’s text embedding-3-small model.
To support user navigation and cluster discovery, the high-dimensional embeddings are reduced via UMAP and grouped using HDBSCAN. A key component of the pipeline is the integration of GPT-based cluster labeling, which generates short, descriptive names for each cluster based on the titles and descriptions of the papers it contains. This step is fully automated but subject to human validation when new clusters emerge; meaning, these clusters can be modified by humans if needed.
4.1.2. Balancing Automation and Oversight.
Unlike other prior approaches that rely on manual curation or entirely unsupervised grouping, our pipeline adopts a semi-automated architecture designed to scale with minimal human intervention while retaining curatorial oversight. Specifically, the system regularly ingests new suggestions, both from automated Semantic Scholar queries and from users interacting with the front-end interface, and proposes candidate entries for including in the pool of benchmarks. Although metadata extraction and embedding are fully automated, curators retain the ability to approve or reject new data sources, review GPT-generated labels, and fine-tune cluster boundaries as needed.
4.1.3. Interactive Visualization.
The final layer of the pipeline provides an interactive 2D interface that enables users to explore benchmarks based on their semantic similarity. The visualization supports fuzzy search, local citation graphs, and interactive paper details. Importantly, user interactions on the frontend—such as suggesting new benchmarks—are routed back into the data collection layer (cf. Figure 5).
4.2. Additional Features
To improve the search and exploration experience, we incorporated several key functionalities. First, a text-based search interface allows users to find articles by title and abstract using fuzzy search via the Fuse.js library [^20] https://www.fusejs.io/. Users can also issue advanced field-specific queries (e.g., language:python) or exact phrase queries (e.g., "code generation"), and benefit from autocomplete suggestions for common terms, programming languages, tasks, and datasets.
A dynamic 2-D UMAP visualization displays papers embedded in semantic space. Hovering reveals abstracts via a Paper Content Tooltip, while double-clicking redirects to the paper’s DOI page. Paper nodes are sized based on normalized citation count, colored by cluster, and annotated with metadata such as year, tasks, and dataset names.
Clicking a point activates the Related Papers feature, which shows (1) The top-5 most similar papers based on cosine similarity, (2) A detailed paper overview with title, abstract, authors, venue/journal, and publication type, (3) A list of citing papers with sortable tables and filters by year and venue, and (4) An interactive Paper Citations Graph, which visualizes inter-citation relationships among citing papers.
Users can navigate between chart, list, and grid views, paginate results, and apply advanced filters by programming/natural language, dataset, task, cluster, year, and citation count. A search history and keyboard shortcuts improve usability. The entire interface supports responsive interaction with detailed feedback and error handling.
Data Collection Metadata & Embeddings Clustering & Labeling Frontend InterfaceUser SuggestionsCuratedGoogle SheetSS API SuggestionsMetadataEnrichmentGitHub / HFDescriptionsOpenAIEmbeddingsUMAPReductionHDBSCANClusteringGPT-basedLabelingInteractive 2D Map+ Fuzzy Search+ Related Papers
Figure 5. Pipeline architecture of BenchScout. This includes automated metadata extraction, semantic embedding, clustering, and interactive visualization.
4.3. User Study
@@
4.3.1. Questionnaire Design
Table 20. Questionnaire Design Overview
@@ 4.3.1. Questionnaire Design
Participant Background What is your professional background?
What is your role?
How familiar are you with AI4SE benchmarks?
How many years of experience do you have in this field?
5-point Likert Scale
(1: Not familiar, 5: Very familiar)
Experience question
(¡1, 1–3, 3–5, 5+ years)
Search Functionality How easy was it to navigate the interface?
How intuitive was the search functionality?
How effective was the tool in finding benchmarks?
Was the visual evaluation of datasets useful?
5-point Likert Scale
(1: Not useful, 5: Extremely useful)
N/A
Cross-referencing Feature How useful was the cross-referencing feature?
Did the tool help in understanding relationships between benchmarks?
Was the visual interface for benchmark similarity useful?
5-point Likert Scale, Open-ended Includes qualitative feedback option
User Experience & Feedback How would you rate the overall user experience?
How likely are you to use the tool in your work?
Did you experience any issues or challenges?
What other tools do you use for searching benchmarks?
How does this tool compare to others?
How well does the tool meet the needs of professionals in AI4SE?
Likert Scale, Open-ended Open-ended questions for in-depth feedback
The questionnaire designed for the study was divided into four key sections to gather feedback about the tool, namely the participant’s background, the quality of the search functionality, the quality of the cross-referencing feature, and the overall user experience. Table 20 is an overview of the posed questions.
4.3.2. Results and Analysis
@@
The respondents of the questionnaire were generally familiar with AI4SE, with an average familiarity rating of 3.8. There were varying levels of experience in the field, with the range between 1-3 years being most common (eight people). More concretely, in our pool of participants, the demographic distribution was as follows:
  1. (1)
    Roles: The participants held diverse roles, including 6 Researchers, 5 PhD Candidates, 5 Students, 4 Software/Research Engineers, and 2 Lead Researchers
  2. (2)
    Experience Level: 8 participants had 1-3 years, 6 had 3-5 years, 3 had 5+ years, and 5 had less than 1 year of experience
  3. (3)
    Familiarity with AI4SE: On a scale of 1 (Not familiar) to 5 (Very familiar), 8 participants rated themselves a ”5” and 7 rated themselves a ”4,” indicating a strongly informed participant pool
@@ 5. BenchFrame
The HumanEval benchmark has long been the de facto standard for evaluating the code-generation capabilities of AI4SE models. It has recently been used to evaluate the latest and greatest LLMs from large companies such as Google Gemini (Gemini Team, 2023) and OpenAI GPT4 (OpenAI, 2023). Despite the great fame, however, our review in Section 3 points towards the existence of numerous notable problems with this foundational benchmark, namely, the existence of incorrect tests, suboptimal canonical solutions, and imprecise problem definitions amongst many others.
5.1. Approach
Full Code Review of HumanEval Standardized Observations Check Observations HumanEval Variants
Modify HumanEval ( $\rightarrow$  HumanEvalNext  ) Peer-Review Changes Evaluate AI4SE Models
Figure 6. Outline of the general approach behind BenchFrame through a case-study of HumanEval.
To improve the quality of the given benchmark, we pursue the following approach – also illustrated in Figure 6. First, we initiate a comprehensive code review, leading to standardized observations (Section 5.1.1). Then, we address these identified issues through a series of modifications (Section 5.1.2), followed by a peer review (Section 5.1.3) to ensure accuracy and reliability. Finally, we experiment with the revised benchmark to evaluate and discuss the results. Below, more details are provided for specific steps in this approach.
5.1.1. Standardized Observations in Current HumanEval Benchmarks
Upon examining and manually reconstructing canonical solutions and experimenting with various HumanEval benchmark test suites, we identified several recurring issues. The system frequently produces incorrect and sub-optimal code, as canonical solutions are inefficient and fail to address critical assumptions outlined in the problem descriptions.
Additionally, these solutions often lack type annotations, further complicating the evaluation process. Another significant problem is the absence of quality testing. Test suites tend to overlap with example tests from the prompt, allowing incorrect canonical solutions to pass. Moreover, there are instances where the expected outputs in the test suites do not align with the canonical solutions’ actual performance.
@@ 5.1.2. Modifications in HumanEvalNext
In HumanEvalNext, we address all the above issues by manually modifying all problems in the original HumanEval benchmark. We decide to improve the original HumanEval as (1) flaws in the original version persist even in improved versions such as HumanEvalPlus, and (2) the fact that HumanEval is still widely used in new literature. While in our we have gone for the original benchmark, due to the adaptable nature of BenchFrame one could opt to have an improved version of the dataset as the starting point. To summarize, these are the general changes made in HumanEvalNext and their benefits: In this work, several key improvements have been made to address the shortcomings of previous benchmarks. First, all suboptimal and incorrect canonical solutions have been fixed, which were previously missed due to insufficient testing and a lack of comprehensive quality review. Furthermore, type annotations have been added to all problems, offering valuable context and simplifying the translation to other programming languages. The original HumanEval benchmark only included type annotations for the first 30 problems, representing merely 18% of the total set of 164 problems. In addition, better support for language conversion frameworks has been incorporated. For instance, HumanEvalNext now features improved compatibility with frameworks like MultiPL-E (Cassano et al., 2022), which supports translation to 18 additional programming languages by standardizing all tests to equality assertions where feasible. This change has reduced the number of incompatible problems by a factor of ten.
@@
Besides these adjustments, challenging scenarios (such as negative values, zero instances, empty inputs, and non-alphanumeric symbols) are incorporated into each task to guarantee that only top-quality AI4SE models, adept at addressing diverse situations, succeed. Accordingly, assertions are implemented within the code wherever constraints are detailed in the problem description. This measure prevents models from ignoring crucial details, thus improving the benchmark’s worth. In particular, we employ specification-based testing to assess functions; using boundary analysis, we explore combinations of within, on, and outside points. Furthermore, the test examples in the problem descriptions have been refined, which significantly affects model performance (Austin et al., 2021). Problems with excessive test examples now feature a reduced set, distributing the evaluation workload more fairly. Lastly, spelling errors have been corrected, descriptions have been consistently formatted, and problem descriptions have been aligned with the implementations, while still leaving room for models to demonstrate intuitive problem-solving skills expected of high-quality AI4SE models. The difficulty level has also been raised by incorporating more edge cases and modifying various problems, aiming to better reflect real-world challenges faced by engineers and to mitigate issues related to data leakage and saturated performance on the HumanEval leaderboards.
To illustrate the modifications of the tests in HumanEvalNext, consider Table 21.
@@
Table 21. Comparison of test statistics between HumanEval (based on https://github.com/openai/human-eval/blob/master/data/HumanEval.jsonl.gzhuman-eval-v2-20210705.json) and HumanEvalNext.
Metric HumanEval (Original) HumanEvalNext $\Delta$
Total number of asserts 1325 2551 $\times 1.92$
Average number of asserts 8 16 $\times 2$
Median number of asserts 7 11 $\times 1.57$
Min. number of asserts 1 4 +3+3
Problems with ¡ 5 asserts 34 2 -94%
5.1.3. Peer Review Process of HumanEvalNext
@@
To ensure the accuracy and reliability of the initial version of HumanEvalNext, an independent reviewer verified all changes. This thorough review involved verifying the clarity and completeness of the problem docstrings, checking for consistency between the problem descriptions and the canonical solutions, and ensuring that both the solutions and test cases were correct and efficient. Where inefficiencies were identified, suggestions for optimization were provided. The review also scrutinized the test cases to ensure comprehensive coverage, identifying any gaps that could allow incorrect solutions to pass. While the initial creation of the benchmark took over 100 hours, the independent peer-review process required an additional 16 hours. As a result of this review, 1% of the problems were redesigned due to structural issues, 9% received additional test cases, and 15% underwent minor grammar or clarity improvements. All suggested changes were documented and reviewed by the original author, with every recommendation either implemented or refined further upon discussion. Since the peer-review involved modifications beyond the test suites, completions for all models were re-run. Despite these changes, the results remained largely consistent with the original, with 40% of models showing no change in their pass@1 scores, 50% showing a 1-2% absolute change, and only 10%—previously top performers—experiencing a 5% drop. This demonstrated that the peer-review process upheld the benchmark’s robustness while refining its quality.
5.2. Experimental Setup
@@
To assess the impact of the modifications applied to the benchmark, we examine the pass@1 performance of ten state-of-the-art open-source software code models using the original HumanEval alongside two enhanced variants, HumanEvalNext and EvalPlus (Liu et al., 2023a). We selected the top-performing models from the big code LLM leaderboards at the evaluation’s start: “NTQAI/Nxcode-CQ-7B-orpo”, “Qwen/CodeQwen1.5-7B”, “deepseek-ai/deepseek-coder-6.7b-instruct”, “TechxGenus/starcoder2-15b-instruct”, “ise-uiuc/Magicoder-S-DS-6.7B”, “Artigenz/Artigenz-Coder-DS-6.7B”, “HuggingFaceH4/starchat2-15b-v0.1”, “google/codegemma-7b-it”, “codeLlama/CodeLlama-13b-Instruct-hf”, and “Stabilityai/stable-code-3b”. For each task in the benchmark (164 total), the LLM is prompted using an instructional preamble asking the model to finish the implementation of the function requested in addition to providing the imports, function header, and function description with each request. We run the inference for the models on a cluster with one NVIDIA A100 80GB GPU and 32 CPU cores. During test execution, a timeout limit of 15 seconds is utilized per function call to disregard completions that are potentially looping forever or are considered overly inefficient with regard to the canonical solutions. Each test is executed in an evaluation suite using the precautions deployed by OpenAI.
5.3. Results